OpenStudy (anonymous):

Which of the following is approximately equal to the quotient of the division problem below? 15/4 ÷ 13/14 a. 4 b. 8 c. 3 d. 6

myininaya (myininaya):

\[a/b \div c/d=a/b \cdot d/c =\frac{a}{b} \cdot \frac{d}{c}=\frac{ad}{bc}\]

myininaya (myininaya):

So flip the second quotient and multiply straight across